OpenStudy (anonymous):

The function f(x)=1/(1+25x^2) is represented as a power series: f(x)=∑n=0∞cnxn Find the first few coefficients in the power series. c0= c1= c2= c3= c4= Find the radius of convergence R of the series. R=

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

for some reason I can't get c1-c5

OpenStudy (anonymous):

I've gotten c0 and R

OpenStudy (anonymous):

Hi :)

OpenStudy (misty1212):

the most well known one is \[\frac{1}{1-x}=\sum x^n\] right?

OpenStudy (anonymous):

Yeah I've got that

OpenStudy (anonymous):

I've even gotten a summation that I thought worked

OpenStudy (misty1212):

replace \(x\) by \(-25x^2\)

OpenStudy (anonymous):

but apparenlty it's wrong

OpenStudy (anonymous):

yup I did that :)

OpenStudy (misty1212):

did you get \[1-25x^2+625x^4-...\]

OpenStudy (anonymous):

I thought the summation is \[\sum_{0}^{\inf} (-25x^2)^n\]

OpenStudy (misty1212):

that is the first three anyways the number get big fast

OpenStudy (anonymous):

Yes that's what I got

OpenStudy (anonymous):

but the program I'm using for my uni says that's wrong

OpenStudy (misty1212):

lets check

OpenStudy (anonymous):

maybe the program is just wrong

OpenStudy (misty1212):

maybe they just want you to write it differently

OpenStudy (misty1212):

\[\sum_{n=0}^{\infty} (-1)^{n}(25x)^{2n}\]

OpenStudy (anonymous):

It just want's the coefficients, I even tried (-1)^1(5)^2

OpenStudy (anonymous):

I think you meant \[\sum_{0}^{\inf}(-1)^n(5x)^{2n}\]

OpenStudy (anonymous):

Thanks at least I feel I'm not going crazy now. I will check with my prof later